# SenseFlow: Scaling Distribution Matching for Flow-based Text-to-Image Distillation

### v1

@@ Abstract
The Distribution Matching Distillation (DMD) has been successfully applied to text-to-image diffusion models such as Stable Diffusion (SD) 1.5. However, vanilla DMD suffers from convergence difficulties on large-scale flow-based text-to-image models, such as SD 3.5 and FLUX. In this paper, we first analyze the issues when applying vanilla DMD on large-scale models. Then, to overcome the scalability challenge, we propose implicit distribution alignment (IDA) to regularize the distance between the generator and fake distribution. Furthermore, we propose intra-segment guidance (ISG) to relocate the timestep importance distribution from the teacher model. With IDA alone, DMD converges for SD 3.5; employing both IDA and ISG, DMD converges for SD 3.5 and FLUX.1 dev. Along with other improvements such as scaled up discriminator models, our final model, dubbed SenseFlow, achieves superior performance in distillation for both diffusion based text-to-image models such as SDXL, and flow-matching models such as SD 3.5 Large and FLUX. The source code will be avaliable at https://github.com/XingtongGe/SenseFlow.
Figure 1: 1024×1024 samples produced by our 4-step generator distilled from FLUX.1-dev.
@@ 1 Introduction
Significant advancements have been made on diffusion models ho2020denoising ; rombach2022high ; podell2023sdxl ; esser2024scaling ; flux2024 for text-to-image generation over recent years. However, these models typically require multiple denoising steps to generate high-quality images. As models continue to scale up in terms of the parameter size, the computational cost and inference time for image generation increase substantially, making the process slower and more resource-intensive. To address this issue, various diffusion distillation methods have been developed to distill a diffusion model into a few-step generator, including consistency models song2023consistency ; luo2023latent ; wang2024phased , progressive distillation salimans2022progressive ; ren2024hyper , adversarial distillation Wang2023DiffusionGAN ; sauer2024adversarial ; chadebec2025flash , and score distillation yin2024one ; yin2024improved ; Luo2023diffinstruct . Currently, the Distribution Matching Distillation (DMD) series yin2024improved have demonstrated superior results in distilling standard diffusion models such as SD 1.5 rombach2022high and SDXL podell2023sdxl .
@@
However, few of these methods have successfully demonstrated effective distillation performance across a broader range of models, particularly in flow-based diffusion models with larger parameter sizes, such as SD3.5 Large (8B) esser2024scaling and FLUX.1 dev (12B) flux2024 . As models increase in architecture complexity, parameter size, and training complexity, it becomes significantly more challenging to distill these models into efficient few-step generators (e.g., a 4-step generator).
@@
In this paper, we introduce SenseFlow, which selects the framework of DMD2 yin2024improved as a touchstone, and scales it up for larger flow-based text-to-image models, including SD3.5 Large and FLUX.1 dev. Specifically, vanilla DMD2 has difficulty in converging and faces significant training instability on large models, even with the time-consuming two time-scale update rule (TTUR) ttur applied. To address this challenge, we propose implicit distribution alignment (IDA) to regularize the distance between the generator and the fake distribution network, which makes the training of fake distribution network faster and easier. This further allows us to make the generator converge more stably.
@@
Further, DMD2 and most existing diffusion distillation methods still use uniformly sampled timesteps for training and inference. However, due to the complex strategies employed during training of teacher diffusion models, different timesteps exert varying denoising effects throughout the entire process, which is also discussed in RayFlow shao2025rayflow . To avoid the inefficiency of naive timestep sampling strategy in distillation, we propose to relocate the teacher’s timestep-wise denoising importance into a small set of selected coarse timesteps. For each coarse timestep $\tau_{i}$, we construct an intra-segment guidance (ISG) by sampling an intermediate timestep $t_{mid}\in(\tau_{i-1},\tau_{i})$, and construct a guidance trajectory: the teacher denoises from $\tau_{i}$ to $t_{mid}$, then the generator continues from $t_{mid}$ to $\tau_{i-1}$. We then guide the generator to align its direct prediction from $\tau_{i}$ to $\tau_{i-1}$ with this trajectory. This guidance mechanism effectively aggregates the teacher’s fine-grained behavior within each segment, improving the generator’s ability to approximate complex transitions across fixed sparse timesteps.
@@
For further enhancement, we incorporate a more general and powerful discriminator built upon vision foundation models (e.g., DINOv2 oquab2023dinov2 , CLIP radford2021learning ), which operates in the image domain and can provide stronger semantic guidance. The use of pretrained vision backbones introduces rich semantic priors, enabling the discriminator to better capture image-level quality and fine-grained structures. By aggregating timestep-aware adversarial signals, this design yields stable and efficient training with superior visual qualities.
@@
  • We discover that vanilla DMD2 suffers from the convergence issue on large-scale text-to-image models, even with TTUR introduced. To tackle this challenge, we propose implicit distribution alignment to regularize the distance between the generator and fake distribution.
@@
  • Experimental results show that our final model, dubbed SenseFlow, achieves superior performance in distilling large-scale flow-matching models ( e.g., SD 3.5, FLUX.1 dev) and diffusion-based models (e.g., SDXL). Our SD 3.5 Based-SenseFlow achieves state-of-the-art 4-step generation performance among all open-source models evaluated in our study.
@@ 2.1 Diffusion Models
[…]
@@ 2.2 Distribution Matching Distillation
From now on we assume a pre-trained diffusion model is available, with learned score function $s_{r}(X_{t},t)$ and distribution $p_{r}(X_{t})$. The Distribution Matching Distillation (DMD) yin2024one ; yin2024improved distills the diffusion model by a technique named score distillation poole2022dreamfusion . More specifically, DMD learns the generator distribution $p_{g}(X_{t})$ to match the diffusion distribution $p_{r}(X_{t})$:
@@
[…]
In practice, the fake distribution is parameterized as the score function $s_{\phi}(X_{t},t)=\nabla\log p_{f}(X_{t})$. On the other hand, the generator is parameterized with a clean image generating network $G_{\theta}(\epsilon),\epsilon\sim\mathcal{N}(0,I)$ and forward diffusion process $q(X_{t}|X_{0})$, such that $p_{g}(X_{t})=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}[q(X_{t}|G_{\theta}(\epsilon))]$. To this end, the DMD updates are achieved by gradient descent and score matching vincent2011connection :
[…]
@@ 3 Method: Scaling Distribution Matching for General Distillation
Figure 2: Left: The generator $\mathcal{G}$ receives a text prompt and $x_{\tau_{i}}$ to produce one-step output $x_{g}$, which is diffused to $x_{t}$ and processed by $s_{\phi}$ and $s_{r}$ for computing DMD gradient. ISG guides $\mathcal{G}$ using an intermediate point $x_{t_{mid}}$, and IDA aligns $\mathcal{G}$ with $s_{\phi}$ after generator update. Right: The discriminator extracts semantic features from generated and real images using CLIP and DINOv2, which are processed by head blocks ${h_{\theta_{i}}}$ to predict real/fake logits for adversarial training. Trainable modules are shown in pink, while frozen (pretrained) ones are shown in grey.
3.1 Bottlenecks in Vanilla DMD series: Stability, sampling, and naive discriminator
@@
3.2 Implicit Distribution Alignment via Generator-Fake Distribution Fusion
In Distribution Matching Distillation (DMD), a critical challenge lies in stabilizing the fake distribution model to accurately track the generator distribution $p_{g}$, especially when working with modern large-scale diffusion backbones such as SD3.5 or FLUX. As model capacity increases and the training strategies of teacher models vary across architectures, ensuring a well-trained fake distribution model becomes increasingly difficult. For example, many models diffusiondpo ; playgroundv3 ; seedream2 use complex post-training strategies to improve the performance of the model in specific directions, such as text rendering or aesthetic quality, which may introduce non-uniform sampling trajectories, making the standard diffusion loss less effective for supervising fake distribution model training.
@@ 3.2 Implicit Distribution Alignment via Generator-Fake Distribution Fusion
Figure 3: “Training Hours-FID” curves on COCO-5K dataset. IDA improves training stability across TTUR ratios.
To address this issue, DMD2 used the two time-scale update rule (TTUR), which increases the update frequency of the fake distribution model relative to the generator. However, TTUR becomes increasingly expensive and brittle as the model size scales up. Results in Fig. 3 also indicate that sometimes even a high ratio of 20:1 still cannot stabilize the training.
On the other hand, although the generator and fake distribution network are optimized via different objectives, their long-term goals are highly aligned: both aim to model a distribution $p_{g}(X_{t})$ that closely approximates the real data distribution $p_{r}(X_{t})$. In practice, they are initialized from the same pretrained teacher and both define the generator-induced distribution $p_{g}(X_{t})$. The key difference is that, generator is guided by an explicit, fixed teacher score $s_{r}(X_{t},t)$ through the variational gradient in Eq. 7 and thus evolves in a clear direction toward $p_{r}$. In contrast, the fake distribution network is trained to regress toward the score of the generator-induced distribution via $\mathcal{L}_{f}$ in Eq. 7, where the target $\nabla\log p_{g}(X_{t})$ is approximated through the generator’s outputs. In early training, this target is a rapidly moving and highly unreliable signal—making the fake distribution network prone to underfitting, drift, or misaligned gradients, especially when the model size is relatively large.
We address this challenge by introducing Implicit Distribution Alignment (IDA), a simple yet effective stabilization mechanism. Specifically, after each generator update, we partially align the fake distribution parameters toward the generator:
$\phi\leftarrow\lambda\cdot\phi+(1-\lambda)\cdot\theta.$ (8)
Intuitively, this allows us to propagate the teacher’s stable supervision—received by the generator—into the fake distribution model indirectly. Since both networks share initialization and long-term alignment, IDA can implicitly regularize the distributional distance between the fake distribution model and generator, preventing it from being misled by the drift moving targets during early training.
@@
In practice, this strategy ensures that the fake distribution remains closely aligned with the generator’s distributional trajectory, especially early in training when score updates are unstable. We observe that combining IDA with even a relatively small TTUR (e.g., 5:1) leads to significantly more stable convergence. An example of this effect is shown in Fig. 3, where we compare FID curves under different TTUR settings with and without IDA. As the figure illustrates, IDA consistently reduces FID variance and improves overall performance. We leave a detailed analysis to the ablation study section.
3.3 Generator Turn: Relocate the Timestep Importance Distribution
Figure 4: Left: The normalized reconstruction errors over timesteps in $[0,1]$. Right: An illustration of the Intra-Segment Guidance.
On the other hand, the distillation performance of vanilla DMD2 is fundamentally limited by fixed timestep supervision. In vanilla DMD2 setups, the generator is only trained at a small set of pre-defined timesteps (e.g., $\tau\in\{249,499,749,999\}$). However, this fixed design introduces two major issues: first, the generator receives no training signal from the rest of the trajectory, which leads to poor generalization for the full trajectory; second, the effectiveness of each supervised timestep is highly sensitive to where it lies along the trajectory—neighboring timesteps can exhibit drastically different predictive errors. To better illustrate the local reliability of different timesteps in the diffusion trajectory, we visualize the normalized one-step reconstruction loss $\xi(t)$ over 1000 uniformly spaced timesteps in $[0,1]$:
$\xi(t):=\mathbb{E}_{x_{0},\epsilon\sim\mathcal{N}(0,I)}\left[\|\hat{x}_{0}(x_{t},t)-x_{0}\|^{2}\right],$ (9)
where $x_{0}$ is generated by the teacher model (SD 3.5 or FLUX.1 dev) and $x_{t}$ is obtained via diffusion forward process in Eq. 4 using $x_{0}$ and $\epsilon$. The results are shown in Fig. 4 Left. We observe that as $t$ increases, the denoising error $\xi(t)$ does not grow monotonically, but instead exhibits noticeable local oscillations—particularly in the interval $t\in[0.8,1.0]$. This suggests that even adjacent timesteps within the same region may differ significantly in their denoising accuracy, implying that their relative “importance” to the overall denoising process is not uniform. Consequently, selecting supervision points without considering their local reliability may inadvertently anchor the generator to suboptimal points, degrading sample quality and training stability.
To mitigate this issue, we propose to relocate the teacher’s denoising importance into a small set of selected coarse timesteps. For each coarse timestep $\tau_{i}$, we construct an intra-segment guidance by randomly sampling an intermediate timestep $t_{1}\in(\tau_{i-1},\tau_{i})$. As shown in Fig. 4 Right, the teacher model generates $x_{t_{1}}$ by denoising from $\tau_{i}$ to $t_{1}$. Then, the generator continues the denoising process from $t_{1}$ to $\tau_{i-1}$, yielding the guidance target $x_{\text{tar}}$. Meanwhile, the generator also produces $x_{\tau_{i-1}}$ directly from $\tau_{i}$ to $\tau_{i-1}$. We then apply an $\mathcal{L}_{2}$ loss between $x_{g}$ and $x_{\text{tar}}$, where gradients are only propagated through the generator path:

### v2

# arXiv:2506.00523v2[cs.CV] 28 Feb 2026

@@ Abstract
The Distribution Matching Distillation (DMD) has been successfully applied to text-to-image diffusion models such as Stable Diffusion (SD) 1.5. However, vanilla DMD suffers from convergence difficulties on large-scale flow-based text-to-image models, such as SD 3.5 and FLUX. In this paper, we first analyze the issues when applying vanilla DMD on large-scale models. Then, to overcome the scalability challenge, we propose implicit distribution alignment (IDA) to constrain the divergence between the generator and the fake distribution. Furthermore, we propose intra-segment guidance (ISG) to relocate the timestep denoising importance from the teacher model. With IDA alone, DMD converges for SD 3.5; employing both IDA and ISG, DMD converges for SD 3.5 and FLUX.1 dev. Together with a scaled VFM-based discriminator, our final model, dubbed SenseFlow, achieves superior performance in distillation for both diffusion based text-to-image models such as SDXL, and flow-matching models such as SD 3.5 Large and FLUX.1 dev. The source code is available at https://github.com/XingtongGe/SenseFlow.
Figure 1: 1024×1024 samples produced by our 4-step generator distilled from FLUX.1-dev. Please zoom in for details.
@@ 1 Introduction
Significant advancements have been made on diffusion models (Ho et al., 2020; Rombach et al., 2022; Podell et al., 2024; Esser et al., 2024; Labs, 2024) for text-to-image generation over recent years. However, these models typically require multiple denoising steps to generate high-quality images. As models continue to scale up in terms of the parameter size, the computational cost and inference time for image generation increase substantially, making the process slower and more resource-intensive. To address this issue, various diffusion distillation methods have been developed to distill a diffusion model into a few-step generator, including consistency models (Song et al., 2023; Luo et al., 2023a; Wang et al., 2024a), progressive distillation (Salimans & Ho, 2022; Ren et al., 2024), adversarial distillation (Wang et al., 2023a; Sauer et al., 2024b; Chadebec et al., ), and score distillation (Yin et al., 2024b; a; Luo et al., 2023b). Currently, the Distribution Matching Distillation (DMD) series (Yin et al., 2024a) have demonstrated superior results in distilling standard diffusion models such as SD 1.5 (Rombach et al., 2022) and SDXL (Podell et al., 2024).
@@
However, few of these methods have successfully demonstrated effective distillation performance across a broader range of models, particularly in flow-based diffusion models with larger parameter sizes, such as SD3.5 Large (8B) (Esser et al., 2024) and FLUX.1 dev (12B) (Labs, 2024). As models increase in architecture complexity, parameter size, and training complexity, it becomes significantly more challenging to distill these models into efficient few-step generators (e.g., a 4-step generator).
@@
In this paper, we introduce SenseFlow, which selects the framework of DMD2 (Yin et al., 2024a) as a touchstone, and scales it up for larger flow-based text-to-image models, including SD3.5 Large and FLUX.1 dev. Specifically, vanilla DMD2 has difficulty in converging and faces significant training instability on large models, even with the time-consuming two time-scale update rule (TTUR) (Heusel et al., 2017) applied. Viewing DMD as a min–max game, the inner best response requires the fake distribution model to track and predict the data distribution determined by generator samples effectively, which is brittle and expensive to realize. We therefore introduce implicit distribution alignment (IDA): a lightweight proximal update applied after each generator step that nudges the fake model toward the generator, maintaining $p_{f}(x_{t})\approx p_{g}(x_{t})$—an $\varepsilon$-best response. This simple alignment markedly improves stability and enables convergence on large backbones.
@@
Further, DMD2 and most existing diffusion distillation methods still use uniformly of handcrafted sampled timesteps for training and inference. However, due to the complex strategies employed during training of teacher diffusion models, different timesteps exert varying denoising effects throughout the entire process, which is also discussed in RayFlow (Shao et al., 2025). To avoid the inefficiency of naive timestep sampling strategy in distillation, we propose to relocate the teacher’s timestep-wise denoising importance into a small set of selected coarse timesteps. For each coarse timestep $\tau_{i}$, we construct an intra-segment guidance (ISG) by sampling an intermediate timestep $t_{mid}\in(\tau_{i-1},\tau_{i})$, and construct a guidance trajectory: the teacher denoises from $\tau_{i}$ to $t_{mid}$, then the generator continues from $t_{mid}$ to $\tau_{i-1}$. We then guide the generator to align its direct prediction from $\tau_{i}$ to $\tau_{i-1}$ with this trajectory. This guidance mechanism effectively aggregates the teacher’s fine-grained behavior within each segment, improving the generator’s ability to approximate complex transitions across fixed sparse timesteps.
@@
For further enhancement, we incorporate a more general and powerful discriminator built upon vision foundation models (e.g., DINOv2 (Oquab et al., 2023), CLIP (Radford et al., 2021)), which operates in the image domain and can provide stronger semantic guidance. The use of pretrained vision backbones introduces rich semantic priors, enabling the discriminator to better capture image-level quality and fine-grained structures. By aggregating timestep-aware adversarial signals, this design yields stable and efficient training with superior visual qualities.
@@
  • We discover that vanilla DMD2 suffers from the convergence issue on large-scale text-to-image models, even with TTUR applied. To tackle this challenge, we propose implicit distribution alignment to constrain the divergence between the generator and the fake model.
@@
  • Experimental results show that our final model, dubbed SenseFlow, achieves state-of-the-art performance in distilling large-scale flow-matching models ( e.g., SD 3.5, FLUX.1 dev) and diffusion-based models (e.g., SDXL).
@@ 2.1 Diffusion Models
$\textup{DDPM: }q(X_{t}\mid X_{0})=\mathcal{N}\!\bigl(e^{-\frac{1}{2}\int_{0}^{t}\beta_{s}ds}X_{0},\;(1-e^{-\frac{1}{2}\int_{0}^{t}\beta_{s}ds})I\bigr),$ (3)
$\textup{FM-OT: }q(X_{t}\mid X_{0})=\mathcal{N}\!\bigl((1-t)X_{0},\;t^{2}I\bigr).$ (4)
@@ 2.2 Distribution Matching Distillation
From now on we assume a pre-trained diffusion model is available, with learned score function $s_{r}(X_{t},t)$ and distribution $p_{r}(X_{t})$. The Distribution Matching Distillation (DMD) (Yin et al., 2024b; a) distills the diffusion model by a technique named score distillation (Poole et al., 2022). More specifically, DMD learns the generator distribution $p_{g}(X_{t})$ to match the diffusion distribution $p_{r}(X_{t})$:
@@
$\textup{Generator:}\min_{p_{g}}\mathbb{E}_{t\sim[0,T],p_{g}}[\log p_{f}(X_{t})-\log p_{r}(X_{t})],$
$\textup{Fake:}\max_{p_{f}}\mathbb{E}_{t\sim[0,T],p_{g}}[\log p_{f}(X_{t})].$ (6)
In practice, the fake distribution is parameterized as the score function $s_{\phi}(X_{t},t)=\nabla\log p_{f}(X_{t})$. On the other hand, the generator is parameterized with a clean image generating network $G_{\theta}(\epsilon),\epsilon\sim\mathcal{N}(0,I)$ and forward diffusion process $q(X_{t}|X_{0})$, such that $p_{g}(X_{t})=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}[q(X_{t}|G_{\theta}(\epsilon))]$. To this end, the DMD updates are achieved by gradient descent and score matching (Vincent, 2011):
$\textup{Generator: }\nabla_{\theta}\mathcal{L}_{g}=\mathbb{E}_{t\sim[0,T],\epsilon\sim\mathcal{N}(0,I),X_{t}\sim q(X_{t}|G_{\theta}(\epsilon))}[(s_{\phi}(X_{t},t)-s_{r}(X_{t},t))\frac{\partial X_{t}}{\partial\theta}],$
$\textup{Fake:}\nabla_{\phi}\mathcal{L}_{f}=\nabla_{\phi}\mathbb{E}_{t\sim[0,T],\epsilon\sim\mathcal{N}(0,I),X_{t}\sim q(X_{t}|G_{\theta}(\epsilon))}[||s_{\phi}(X_{t},t)-\nabla_{X_{t}}\log q(X_{t}|G_{\theta}(\epsilon))||].$ (7)
@@ 3 Method: Scaling Distribution Matching for General Distillation
Figure 2: Left: The generator $\mathcal{G}$ receives a text prompt and $x_{\tau_{i}}$ to produce one-step output $x_{g}$, which is diffused to $x_{t}$ and processed by $s_{\phi}$ and $s_{r}$ for computing DMD gradient. ISG guides $\mathcal{G}$ using an sampled intermediate $t_{mid}$}, and IDA aligns $\mathcal{G}$ with $s_{\phi}$ after generator update. The overall training pipeline is shown in Algorithm 1. Right: The discriminator extracts semantic features from generated and real images using CLIP and DINOv2, which are processed by head blocks $h_{\theta_{i}}$} to predict real/fake logits for adversarial training. Trainable modules are shown in pink, while frozen (pretrained) ones are shown in grey.
3.1 Bottlenecks in Vanilla DMD: Stability, sampling, and naive discriminator
@@
3.2 Implicit Distribution Alignment (IDA) for Flow-based Models
Recall that the DMD can be viewed as a min-max optimization:
$\min_{\theta}\max_{\phi}V(\theta,\phi)\triangleq\mathbb{E}_{t\sim[0,T],p_{g}}[\log p_{f}(X_{t})-\log p_{r}(X_{t})].$ (8)
It is obvious that the inner best response is attained at $p_{f}(X_{t})=p_{g}(X_{t})$. Furthermore, following Proposition 2 of GAN (Goodfellow et al., 2014), if the inner best response is achieved at every round of generator optimization, then generator converges.
@@ 3.2 Implicit Distribution Alignment (IDA) for Flow-based Models
Figure 3: “Training Hours-FID” curves on COCO-5K dataset. When distilling the 8B SD 3.5 Large, IDA improves training stability across TTUR ratios.
However, the inner best response is brittle. DMD2 uses two time-scale update rule (TTUR) (Heusel et al., 2017), which increases the update frequency of fake model. However, at large-scale models such as SD 3.5 Large (8B), simply increasing TTUR ratio is expensive and can still oscillate, as shown in Fig. 3. To tackle the difficulty to achieve inner best response, we introduce a proximal alignment step after each generator update, called Implicit Distribution Alignment (IDA):
$\phi\leftarrow\lambda\phi+(1-\lambda)\theta,$ (9)
where $\lambda\in(0,1]$ and close to $1$. We claim that IDA maintains $p_{f}(X_{t})\approx p_{g}(X_{t})$, and thus help DMD converges. More formally, we have the following:
Proposition 3.1.
Under mild assumptions (Assumptions A.1 and A.6), IDA maintains an $\epsilon$-best inner response. More specifically, after $k$ round of min max optimization in Eq. 8, we have
$\mathbb{E}_{t}\,D_{KL}\!\big(p_{g}(X_{t})\,\|\,p_{f}(X_{t})\big)\ \leq\ \varepsilon,\textit{ i.e., }p_{f}(X_{t})\approx p_{g}(X_{t})\ \ \text{($\varepsilon$-best response).}$ (10)
@@
In practice, this strategy ensures that the fake distribution remains closely aligned with the generator’s distributional trajectory. We observe that combining IDA with even a relatively small TTUR (e.g., 5:1) leads to significantly more stable convergence. An example of this effect is shown in Fig. 3, where we compare FID curves under different TTUR settings with and without IDA. As the figure illustrates, IDA consistently reduces FID variance and improves overall performance.
Intuitively, the generator is trained on a small set of anchor timesteps (e.g., four fixed $t$’s), while the fake model is trained on a much denser grid. The IDA transfers the generator’s progress on those anchor $t$’s to the fake model, improving its predictions at those $t$’s. On the other hand, on a SD 3.5 Medium generator, we also observe that DMD training produces generators whose samples are consistent across $4/8/16$ Euler steps, indicating that $G_{\theta}$ remains close to a smooth continuous-flow integrator. This observation supports using generator parameters to softly anchor the fake model. Proofs, discussion, and visualization results are provided in Appendix §A.
3.3 Intra-Segment Guidance: Relocating Timestep Importance
The distillation performance of vanilla DMD2 is constrained by supervision at a few handcrafted, coarse timesteps (e.g., $\tau\in\{249,499,749,999\}$). This design has two drawbacks: (i) the generator receives no signal over the rest of the trajectory, hurting generalization; and (ii) the utility of each supervised timestep is highly position-dependent—neighboring timesteps can differ markedly in prediction error. To quantify this local reliability, we visualize the normalized one-step reconstruction loss
$\xi(t):=\mathbb{E}_{x_{0},\,\epsilon\sim\mathcal{N}(0,I)}\!\left[\left\|\hat{x}_{0}(x_{t},t)-x_{0}\right\|_{2}^{2}\right],$ (11)
where $x_{0}$ is sampled from the teacher (SD 3.5 or FLUX.1-dev) and $x_{t}$ is obtained by the forward diffusion process in Eq. 4. As shown in Fig. 4 (Left), $\xi(t)$ is not monotonic in $t$ and exhibits pronounced local oscillations—especially for $t\in[0.8,1.0]$. Hence, treating nearby timesteps as equally informative can anchor training to suboptimal points.
To address this, we introduce Intra-Segment Guidance (ISG), which relocates the teacher’s denoising importance from within each segment $(\tau_{i-1},\tau_{i}]$ to its supervised anchor. For every coarse timestep $\tau_{i}$, we sample an intermediate timestep $t_{\text{mid}}\in(\tau_{i-1},\tau_{i})$. The teacher denoises from $\tau_{i}$ to $t_{\text{mid}}$ to produce $x_{t_{\text{mid}}}$. The generator then continues from $t_{\text{mid}}$ to $\tau_{i-1}$, yielding the target $x_{\text{tar}}$. In parallel, the generator directly denoises from $\tau_{i}$ to $\tau_{i-1}$ to produce $x_{\tau_{i-1}}$. We optimize an $\ell_{2}$ loss that backpropagates only through the generator path: